OpenStudy (anonymous):

y= 3+x What is the y-intercept?

OpenStudy (anonymous):

x=0 then y=...?

OpenStudy (anonymous):

oh, three? Right?

OpenStudy (anonymous):

y intercept is when the function has a y value of zero substitute y=0 and we have 0=3+x x=-3

OpenStudy (anonymous):

right! it's y=3

OpenStudy (anonymous):

THANKS!! :)

OpenStudy (anonymous):

jm92 - you are talking about x-intercept :))

OpenStudy (anonymous):

ok thx for the clarification :]

OpenStudy (anonymous):

do either of you know what the slope is then?

OpenStudy (anonymous):

just think about it as: y-intercept = crossing y-axis (means x=0) x-intercept = crossing x-axis (means y=0)

OpenStudy (anonymous):

if we are talking about original: y= 3+x the slope is the coefficient in a front of x which is m= 1 in this case, right?

OpenStudy (anonymous):

oh, yes!!! THanks Inik, YOU ROCK!

OpenStudy (anonymous):

Good Luck & have fun with math :)))